# Frontiers in Intelligent Colonoscopy

### v1

@@ 5.3.2 Diagnostic studies
$\bullet$ Visual encoder. Our diagnostic experiment begins with an inquiry – What types of visual representations are appropriate for multimodal colonoscopy data? We prepare four sets of representations from various large-scale pre-training strategies: supervised learning (ConvNeXtV2 [285], ViT [100]), reconstructive learning (MAE [286]), and contrastive learning using vision-only data (DINOv2 [287]) or VL data (CLIP [288], SigLIP [5]). As shown in Tab. (8-a), all encoders use pre-trained weights from Huggingface. To ensure consistency, we manually interpolate the smaller position embedding for MAE and DINOv2 from 224px to 384px (marked with *), while leaving the default input of the remaining models unchanged. Our observation reveals that contrastive learning encoders using VL data outperform other strategies. This suggests that visual representations pre-aligned with weak texts during their pre-training facilitate us to transform visual embeddings into the language space. Regarding the other unimodal encoders, both supervised learning methods (ConvNeXtV2, ViT) and the reconstructive approach (MAE) give satisfactory feedback. However, the vision-only contrastive learning model (DINOv2) struggles, especially in the REC and REG tasks. Multiple efforts to optimise DINOv2 with various hyperparameters, including input size, model size, and learning rate, did not produce acceptable results for multimodal colonoscopy tasks.
@@
$\bullet$ Multigranularity multimodal adapter. It serves as a key component in linking vision and language modalities, reducing visual tokens to mitigate computational overhead. As detailed in Tab. (8-b), we analyse its effectiveness from three perspectives. (a) How to configure the pooling kernels? As a reference, we initialise a baseline variant of ColonGPT with a multimodal adapter used in [8, 7], which employs a pure MLP architecture to process all input tokens equally. We then gradually decrease the size of the pooling kernels across five variants: $\{16,8,1\}$, $\{14,7,1\}$, $\{12,6,1\}$, $\{10,5,1\}$, and $\{8,4,1\}$. Concerning the performance-cost trade-offs, our setup $\{14,7,1\}$ is optimal. It decreases the visual tokens from 100% (729 tokens) to 33.74% (246 tokens) while maintaining impressive results across three conversational tasks. To illustrate, we observe performance gains of 2.28%, 1.62%, and 2.07% in the CLS, REG, and REC tasks, respectively. (b) Is global context necessary? We remove the global view from our default setup $\{14,7,1\}$ for the multigranularity adapter, producing a controlled variant with setup $\{14,7\}$. The performance then declines, indicating the necessity of capturing the global context within visual embeddings for improved outcomes. (c) Is positional encoding important? As shown in the last row of Tab. (8-b), our setup without positional encoding shows a significant performance drop in REC task, from 45.77% to 40.09%. This suggests that the relative position information for the visual sequence is crucial for the localisation task.
$\bullet$ Fine-tuning strategy. Lastly, we investigate how to effectively tune our model on multimodal colonoscopy data. As shown in Tab. (8-c), we initiate a set of variants to tune the language model, Phi1.5. It includes seven variants with different setups for LoRA and a full-tuning variant as a reference point. The best performance was observed in the LoRA variant with configuration $r/\alpha=128/256$, achieving 85.81% and 83.42% in the CLS and REG tasks, respectively. In addition, the other two variants achieve outstanding scores in the REC task, with the full-tuning variant achieving 50.20% and the LoRA with configuration $r/\alpha\!=\!256/512$ reaching 48.27%. This implies that more tunable parameters for the language model could allow ColonGPT to capture more complex patterns in the spatial planes. Therefore, we believe that there remains room to improve the localisation ability of ColonGPT in the REC task, such as increasing the rank factor $r$ and finding a matched scaling factor $\alpha$. For training efficiency, we finally choose $r/\alpha\!=\!128/256$ as our default configuration, as it allows us to reduce the training time from eight to roughly five hours.
5.3.3 Empirical takeaways
@@
$\bullet$ Efficiency drives progress. As discussed above, we take less data to obtain greater performance than other model rivals. This success benefits from the way we build ColonGPT. (a) Colonoscopy data inherently contains redundant information, such as the fact that most mucosal surfaces are similar, as well as camouflaged patterns between benign lesions and their surroundings, as discussed in §2.2. To reduce redundancy, we propose a multigranularity multimodal adapter that selectively samples tokens without compromising performance. For future improvement, we can draw on the wisdom of previous token reduction techniques [292]. (b) As shown in Tab. VII, the Phi1.5 model [6], although lightweight, shows surprising efficiency, even outperforming other 7B-level competitors. This indicates that larger models appear to require more colonoscopy data. Thus, future efforts should prioritise enhancing training and inference efficiency, especially for the medical field, rather than racing with massive computational resources. A promising idea to streamline the MLM framework using an encoder-free solution [293] to interpret visual pixels.
@@
6 Conclusion
We investigate the frontiers in intelligent colonoscopy techniques and their broader implications in the multimodal field. Our structure follows two primary threads. First, we survey the landscape of four colonoscopic scene perception tasks and sort out the key challenges and understudied areas. Second, our survey reveals that multimodal research in colonoscopy is underexplored. To embrace this, we contribute three initiatives to the community: a large-scale multimodal instruction tuning dataset ColonINST, a colonoscopy-specific multimodal language model ColonGPT, and a multimodal benchmark. Importantly, multimodal colonoscopy research is rapidly advancing. Future developments can build on recent breakthroughs, such as executing role-playing tasks with agents [297].
@@ Acknowledgments
This research was supported by NSFC (NO.62476143). We express our sincere gratitude to Yu-Cheng Chou (JHU) and Stephen Gould (ANU) for interesting discussions.

### v2

@@ V-D Diagnostic experiments
$\bullet$ Visual encoder. Our diagnostic experiment begins with an inquiry – What types of visual representations are appropriate for multimodal colonoscopy data? We prepare four sets of representations from various large-scale pre-training strategies: supervised learning (ConvNeXtV2 [285], ViT [100]), reconstructive learning (MAE [286]), and contrastive learning using vision-only data (DINOv2 [287]) or VL data (CLIP [288], SigLIP [5]). As shown in Tab. (VIII-a), all encoders use pre-trained weights from Huggingface. To ensure consistency, we manually interpolate the smaller position embedding for MAE and DINOv2 from 224px to 384px (marked with *), while leaving the default input of the remaining models unchanged. Our observation reveals that contrastive learning encoders using VL data outperform other strategies. This suggests that visual representations pre-aligned with weak texts during their pre-training facilitate us to transform visual embeddings into the language space. Regarding the other unimodal encoders, both supervised learning methods (ConvNeXtV2, ViT) and the reconstructive approach (MAE) give satisfactory feedback. However, the vision-only contrastive learning model (DINOv2) struggles [^1] Notably, multiple efforts to optimise DINOv2 with various hyperparameters, including input size, model size, and learning rate, did not produce acceptable results for multimodal colonoscopy tasks., suggesting that its visual representations may be difficult to align with the language space.
@@
$\bullet$ Multigranularity multimodal adapter. It serves as a key component in linking vision and language modalities, reducing visual tokens to mitigate computational overhead. As detailed in Tab. (VIII-b), we analyse its effectiveness from three perspectives. (a) How to configure the pooling kernels? As a reference, we initialise a baseline variant of ColonGPT with a multimodal adapter used in [8, 7], which employs a pure MLP architecture to process all input tokens equally. We then gradually decrease the size of the pooling kernels across five variants: $\{16,8,1\}$, $\{14,7,1\}$, $\{12,6,1\}$, $\{10,5,1\}$, and $\{8,4,1\}$. Concerning the performance-cost trade-offs, our setup $\{14,7,1\}$ is optimal. It decreases the visual tokens from 100% (729 tokens) to 33.74% (246 tokens) while maintaining impressive results across three conversational tasks. To illustrate, we observe performance gains of 5.14% and 4.91% in the CLS and REG tasks, respectively. (b) Is global context necessary? We remove the global view from our default setup $\{14,7,1\}$ for the multigranularity adapter, producing a controlled variant with setup $\{14,7\}$. The performance then declines, indicating the necessity of capturing the global context within visual embeddings for improved outcomes. (c) Is positional encoding important? As shown in the last row of Tab. (VIII-b), our model without positional encoding shows a significant performance drop in the REG task, from 80.18% to 76.35%. This suggests that the relative position information for the visual sequence is crucial for the localisation task.
$\bullet$ Fine-tuning strategy. Lastly, we investigate how to effectively tune our model on multimodal colonoscopy data. As shown in Tab. (VIII-c), we initiate a set of variants to tune the language model, Phi1.5. It includes seven variants with different LoRA setups and a full-tuning variant as a reference point. The best performance was observed in the LoRA variant with configuration $r/\alpha=128/256$. Moreover, the full-tuning variant yields unsatisfactory performance on three tasks. This suggests that an increase in the tunable parameters does not produce gains under the conditions of limited-scale data.
V-E Empirical takeaways
@@
$\bullet$ Efficiency drives progress. As discussed above, we take less data to obtain greater performance than other model rivals. This success benefits from the way we build ColonGPT. (a) Colonoscopy data inherently contains redundant information, such as the fact that most mucosal surfaces are similar, as well as camouflaged patterns between benign lesions and their surroundings, as discussed in §II-B. To reduce redundancy, we propose a multigranularity multimodal adapter that selectively samples tokens without compromising performance. For future improvement, we can draw on the wisdom of previous token reduction techniques [292]. (b) As shown in Tab. VII, the Phi1.5 model [6], although lightweight, shows surprising efficiency, even outperforming other 7B-level competitors. This indicates that larger models appear to require more colonoscopy data. Thus, future efforts should prioritise enhancing parameter efficiency, especially for the medical field, rather than racing with massive computational resources. A promising idea to streamline the MLM framework using an encoder-free solution [293] to interpret visual pixels.
@@
VI Conclusion
We investigate the frontiers in intelligent colonoscopy techniques and their broader implications in the multimodal field. Our structure follows two primary threads. First, we survey the landscape of four colonoscopic scene perception tasks and sort out the key challenges and understudied areas. Second, our survey reveals that multimodal research in colonoscopy is underexplored. To this end, we contribute three initiatives to the community: a large-scale multimodal instruction tuning dataset ColonINST, a colonoscopy-specific multimodal language model ColonGPT, and a multimodal benchmark.
@@ Acknowledgments
This work was supported by the ANU-Optus BRCoE (scholarship awarded to Ge-Peng Ji) and the NSFC (No. 62476143). We express our sincere gratitude to Yu-Cheng Chou (JHU) and Stephen Gould (ANU) for their interesting discussions.